OpenStudy (anonymous):

x / |1-x^2| > 1 / 2 solve the inequalities. someone??

jhonyy9 (jhonyy9):

x different always the values of -/+1

OpenStudy (anonymous):

x<1

OpenStudy (anonymous):

the answer is [ 0.41, 1) U (1, 2.41] how come??

OpenStudy (anonymous):

@jhonyy9 @kumarsajan

OpenStudy (anonymous):

\[\frac{x}{|1-x^2|}>\frac{1}{2}\] \[2x>|1-x^2|\] now \[|1-x^2|=1-x^2 \]on the intervals where \(1-x^2>0\) i.e. if \(x<-1\) or \(x>1\)

OpenStudy (anonymous):

where the number line?? must use or not?

OpenStudy (anonymous):

ok this might be somewhat clearer, maybe not. lets start with \(2x>|1-x^2|\) or maybe \[|(1-x)(1+x)|<2x\] now this can't happen if x is negative, because the absolute value is always positive, so we have to only check on two intervals, \((0,1)\) and \((1,\infty)\)

OpenStudy (anonymous):

okok. i'll try.

OpenStudy (anonymous):

ok the. i understand that x can't be negative. so than. we need to put it up. not trying to make it less then or equal to zero. right?

OpenStudy (anonymous):

well i made a mistake. if \(0<x<1\) we have \(|1-x^2|=1-x^2\) so solve \[1-x^2<2x\] get \[x^2+2x-1>0\] \[x>\sqrt{2}-1\]

OpenStudy (anonymous):

right, and you have to solve two inequalities depending on whether x is in the interval \((0,1)\) or \((1,\infty)\)

OpenStudy (anonymous):

on \((1,\infty)\) you have \(|1-x^2|=x^2-1\) so you have to solve \[x^2-1<2x\] or \[x^2-2x-1<0\]